OpenStudy (sagewilson):

Will Medal. Find the area of a trapezoid if the altitude is 6" and the median is 8." 24 sq. units 48 sq. units 96 sq. units

OpenStudy (sagewilson):

@inkyvoyd

OpenStudy (welshfella):

area of a trapezoid = (a +b) * h/2 where h = altitude and (a + b) / 2 is the median The median is the sum of the parallel sides divided by 2

OpenStudy (sagewilson):

Okay and if I were to draw it out it would look like this right? 

OpenStudy (welshfella):



OpenStudy (welshfella):

M is the median

OpenStudy (welshfella):

area = 6 * 8

OpenStudy (sagewilson):

That would be 48

OpenStudy (welshfella):

yea

OpenStudy (sagewilson):

That's really all I had to do?

OpenStudy (welshfella):

thats correct

OpenStudy (sagewilson):

well, thanks for your help then :)

OpenStudy (welshfella):

yw

OpenStudy (sagewilson):

do you think you can help me with one more?

OpenStudy (welshfella):

yes but plz make another post

OpenStudy (sagewilson):

I will.